Refer to the bond details in Problem 14-1A, except assume that the bonds are issued at a price of $4,895,980. Problem 14-2A SHelp with 1-5 Please
Answer #1
Journal entry
Date General Journal Debit Credit
1/1/2019 Cash 4,895,980
premium on bonds 895,980
bonds payable 4,000,000
2-a) par maturity value Annual rate / year semi annual cash payment
4,000,000 * 6% 6./12 120000
semi annual Straight line
2-b) bond price par value premium periods premium amortization
4,895,980 - 4,000,000 = 895,980 / 30 = 29866
2-c) Semi annual cash premium bond interest expense
payment amortization
120,000 - 29866 = 90,134
3) total bond interest expense over life of bonds
amount repaid
30 payments of 120,000 3600000
par value ant maturity 4,000,000
total repaid 7600000
less amount borrowed 4,895,980
total bond interest expense. 2,704,020
4) unamort Carrying
period premium value
1/1/2019 895,980 4,895,980
06/30/19 866,114 4,866,114
12/31/19 836,248 4,836,248
06/30/20 806,382 4,806,382
12/31/20 776,516 4,776,516
5)
Date General Journal Debit Credit
6/30/19 interest expense 90,134
premium on bonds payable 29,866
cash 120,000
31/12/2019
interest expense 90,134
premium on bonds payable 29,866
cash 120,000